OpenStudy (chris215):

find one of the two positive integers whose product is 196 and whose sum is a minimum. i got 14

OpenStudy (anonymous):

probably the square root right?

OpenStudy (anonymous):

yeah \(\sqrt{196}=14\)

OpenStudy (chris215):

yeah thanks :)

OpenStudy (anonymous):

is it clear, without calculus, that it has to be the square root?

OpenStudy (chris215):

no

OpenStudy (anonymous):

ok you say one number is \(x\) and the other \(y\) and you want to minimize \(x+y\) given that \(xy=196\)

OpenStudy (anonymous):

then you probably changed it to \[y=\frac{196}{x}\] and went on to find the minimum of \[f(x)=x+\frac{196}{x}\] using cacl

OpenStudy (anonymous):

but i say, no it is not "find the minimum of \(x+y\) given \(xy=196\) it is find the minimum of \(y+x\) given \(yx=196\)"

OpenStudy (anonymous):

you say "you idiot, that is exactly the same question doe" and i say yeah right, you can't tell \(x\) from \(y\) one is not more special than the other so the minimum must occur when they are equ8al because you can't tell them apart

OpenStudy (anonymous):

if \(x=y\) and \(xy=196\) then \(x=y=\sqrt{196}\)

OpenStudy (chris215):

ok yeah im pretty sure its 14

OpenStudy (anonymous):

it is 14

OpenStudy (chris215):

thank you so much!